OpenStudy (anonymous):

Simplify: 4+3i / 5-2i

OpenStudy (anonymous):

\[\frac{ 4+3i }{ 5-2i }\]

OpenStudy (anonymous):

when simplifying a complex number in the denominator you generally multiply the fraction by a special case of 1, that is the complex conjugate over the complex conjugate \[ \frac{1}{a+bi}\frac{a-bi}{a-bi}=\frac{a-bi}{a^2 -b^2i^2}=\frac{1-bi}{a^2+b^2}\]

OpenStudy (anonymous):

Thnx!